OpenStudy (anonymous):

umber company can convert logs either lumber or plywood. in a given day the mill will turn out twice as many units of plywood as lumber.it makes a profit 30 on a unit of lumber and 50 on a unit of plywood how many of each must be produced in order to make a profit of 13130

OpenStudy (anonymous):

try making an equation first and set it all equal to =13130

OpenStudy (dumbcow):

2 equations, 2 variables P =2L 30L +50P = 13130

OpenStudy (anonymous):

30 profit on unit of lumber 50 profit on unity of plywood 50(2p)+30l=13130 50(2x)+30x=13130 50*2x=100x 100x+30x=13130 100x+30x=130x 130x=13130 13130/130=101 x=101 101 units must be produced of each to make 13130

OpenStudy (anonymous):

im not sure i understand how do these add up to 13130

OpenStudy (anonymous):

above i made an equation 50(2p)+30l=13130 same as 50(2x)+30x=13130 just put in x for p and l and solve for x...